# The ELM Neuron: an Efficient and Expressive

### v1

@@ 2.1 Architectural Components
Spiking neurons emit their output spike at the axon hillock when the membrane voltage crosses an (approximate and varying) threshold Kandel et al. (2000). For the ELM neuron, we chose a flexible linear readout layer ${\color[rgb]{0.80078125,0.46875,0.73828125}\boldsymbol{w_{y}}}$ on top of the memory units (none of which have to correspond to membrane voltage a priori). The output of the readout layer can be a real value between 0 and 1, representing an output spike probability (in neuroscience attributed as a momentary rate), which ensures higher compatibility with typical machine learning tasks. Therefore, similar to biophysical models (e.g. Hodgkin Huxley Hodgkin and Huxley (1952)), there is no explicit reset mechanism, refractoriness, post-spike dynamics. All of these are implicitly learned by the ELM neuron. To instantiate an ELM neuron as a spiking neural network, a simple output rectification could be applied to recover binary output Zenke and Vogels (2021).
Additional design choices.
We found that enabling greater changes ${\color[rgb]{0.87109375,0.5625,0.01953125}\boldsymbol{\Delta m_{t}}}$ by introducing a multiplicative factor ${\color[rgb]{0.87109375,0.5625,0.01953125}\lambda}$ (a hyperparameter) improved training performance. Furthermore, note that the factor ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\kappa_{m}}}$ is parametrized with $\exp(\nicefrac{{-\Delta t}}{{{\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\tau_{m}}}}})$ (derived from classic neuron models Dayan and Abbott (2005)) and we learn the timescales ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\tau_{m}}}$ directly (constrained using a $sigmoid$ to the the desired range). As opposed to learning ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\kappa_{m}}}$ and initializing it with $1-\Delta t/{\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\tau_{m}}}$, we get valid decay constants even for ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\tau_{m}}}<\Delta t$, and observed improved training stability.
@@ 3 Related Work
Accurately replicating the full spike-level input/output (I/O) relationship of detailed biophysical neuron models at millisecond resolution in a computationally efficient manner presents a formidable challenge. However, addressing this task could potentially yield valuable insights into neural mechanisms of expressivity, learning, and memory capabilities.
@@
The relative scarcity of prior work on this subject can be partially attributed to the computational complexity of cortical neurons only recently garnering increased attention Tzilivaki et al. (2019); Beniaguev et al. (2021); Larkum (2022); Poirazi and Papoutsi (2020). Additionally, traditional phenomenological neuron models have primarily aimed to replicate specific computational phenomena of neurons or networks Koch (1997); Izhikevich (2004); Herz et al. (2006), rather than the entire I/O relationship. Finally, the computational resources and tooling necessary for an extensive exploration of suitable designs have only recently become widely available.
@@
Phenomenological modeling research on temporally or spatially less detailed I/O relationship of biophysical neurons has been primarily centered around the use of multi-layered ANN structures in analogy to the neurons dendritic tree Poirazi et al. (2003); Tzilivaki et al. (2019); Ujfalussy et al. (2018). Similarly, we parametrize the synaptic integration with an MLP, while crucially extending this modeling perspective in several ways. Drawing upon the principles of classical phenomenological modeling via differential equations Izhikevich (2004); Dayan and Abbott (2005), our approach embraces the recurrent nature inherent to neurons. We further consider the significance of hidden states ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m}}$ beyond membrane voltage, as seen in prior works with predetermined variables Brette and Gerstner (2005); Gerstner et al. (2014). This addition enables us to flexibly investigate internal memory timescales ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\tau_{m}}}$, hinted at in earlier modeling studies Gjorgjieva et al. (2016); Cavanagh et al. (2020).
The relationship between timescales and the computations in the brain is the topic of numerous recent investigations in neuroscience. It was shown that intrinsic timescales are diverse across the cortex Murray et al. (2014), which can lead to increased neural expressivity Gjorgjieva et al. (2016); Cavanagh et al. (2020). A common way they occur in computational neuroscience is through the variables capturing slowly building up or leaking out biophysical quantities Hodgkin and Huxley (1952); Mozer (1991); Dayan and Abbott (2005). While timescales can be shaped by task requirements Gao et al. (2020); Zeraati et al. (2023); Cramer et al. (2020a), longer neuronal timescales might be used for temporal credit assignment across larger time horizons Bellec et al. (2018).
@@
Deep learning architectures explicitly incorporating timescales have recently become more prevalent, with several recent architectures based on RNN, transformers or state-space models successfully incorporating them for improved temporal information processing Gu et al. ; Mahto et al. (2021); Smith et al. (2023); Ma et al. (2023). Initially, they appeared in Leaky-RNN Mozer (1991); Jaeger (2002); Kusupati et al. (2018); Tallec and Ollivier (2018), which use a convex combination of old memory and update (as done in ELM using ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\kappa_{m}}}$). In contrast to a time-varying memory decay or complex multi-staged gating architectures, the ELM features a much simpler recurrent cell architecture only using fixed but trainable time constants ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\tau_{m}}}$ and a single MLP to parametrize its integration dynamics.
Comparison with common RNN structures like LSTM Hochreiter and Schmidhuber (1997) and GRU Cho et al. (2014) surface two major differences. First, we set ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\kappa_{m}}}$ as a trainable constant instead of using parametric input and forget gating mechanisms Tallec and Ollivier (2018). Second, we utilize an MLP instead of a simpler parametric affine transformation for calculating ${\color[rgb]{0.87109375,0.5625,0.01953125}\boldsymbol{\Delta m_{t}}}$. Therefore, in contrast with most previous research, our architecture puts the primary importance not on the various gates, also leaving the forget gate constant during inference (which is otherwise believed to be hugely important Van Der Westhuizen and Lasenby (2018)), but rather on the memory update vector calculation itself.
@@ 4 Experiments
In the experimental section of this work, we address three primary research questions. First, are the inductive biases imparted in the ELM neuron adequate to accurately fit a high-fidelity biophysical simulation with a small number of parameters? We detail this investigation in Section 4.1. Second, is the ELM neuron capable of effectively integrating temporal information? We explore this issue in Section 4.2. Third, we test whether the ELM neuron is expressive enough to solve tasks with complex and long-range dependencies, which is discussed in Section 4.3. Each of these questions contributes to our overall understanding of the ELM neuron’s capabilities and its potential applications in both biological and machine learning contexts.
4.1 Fitting a complex biophysical cortical neuron model’s input-output relationship
@@
The NeuronIO dataset primarily consists of simulated input-output (I/O) data for a complex biophysical layer 5 cortical pyramidal neuron model Hay et al. (2011). Input data features biologically inspired spiking patterns (1278 pre-synaptic spike channels), while output data comprises the model’s somatic membrane voltage and output spikes (see Figure 2a and 2b). The dataset and related code is publicly available Beniaguev et al. (2021). The dataset was pre-processed in accordance with Beniaguev et al. (2021), and the models were trained using Binary Cross Entropy (BCE) for spike prediction and Mean Squared Error (MSE) for somatic voltage prediction, with equal weighting.
@@
Figure 2: The ELM neuron is a computationally efficient model of cortical neuron. a) detailed biophysical model of a layer 5 cortical pyramidal cell was used to generate the NeuronIO dataset consisting of input spikes and output spikes and voltage. b) Example target and predictions by the first ELM neuron that crosses the performance threshold. c) and d) Voltage and spike prediction performance of the respective surrogate models. While previous works required around 10M parameters to make accurate spike predictions using a TCN Beniaguev et al. (2021), an LSTM baseline is able to do it with 266K, and our ELM neuron model requires merely 53K, simultaneously achieving much better voltage prediction performance than the TCN. Additional figures comparing FLOPS in appendix.
Figure 3: The ELM neuron gives relevant neuroscientific insights. Ablations on NeuronIO of different hyperparameters of an ELM neuron with AUC $\approx 0.992$, and a (branch) ELM with the same default hyperparameters. Number of removed divergent runs marked with $1^{*}$. a) We find between 10 and 20 memory-like hidden states to be required for accurate predictions, much more than typical phenomenological models use Izhikevich (2004); Dayan and Abbott (2005). b) Highly nonlinear integration of synaptic input is required, in line with recent neuroscientific findings Stuart and Spruston (2015); Jones and Kording (2022); Larkum (2022). c) Allowing greater updates to the memory units is beneficial. d-f) Ablations of memory timescale (initialization and bounding) range or (constant) value, with default range being 1ms-150ms. Timescales around 25 ms seem to be the most useful (matching the typical membrane timescale in the cortex Dayan and Abbott (2005)); however, a lack can be partially compensated by longer timescales. g) and h) Ablating the number of branches $d_{\mathrm{tree}}$ and number of synapses per branch $d_{\mathrm{brch}}$ of the (branch) ELM neuron.
@@
Our ELM neuron achieves better prediction of voltage and spikes than previously used architectures for any given number of trainable parameters (and compute). In particular, it crosses the “sufficiently good” spike prediction performance threshold (0.991 AUC) as defined in (Beniaguev et al., 2021) by using merely 50K trainable parameters, which is around 200$\times$ improvement compared to the previous attempt (TCN) that required around 10M trainable parameters, and 6$\times$ improvement over a LSTM baseline which requires around 266K parameters (see Figure 2c-d). Overall, this result indicates that recurrent computation is indeed an appropriate inductive bias for modeling cortical neurons.
@@
We use the fitted model to investigate how many memory units and which timescales are needed to match the neuron closely. We find that around 20 memory units are required (Figure 3a) with timescales that are allowed to reach at least 25 ms (Figure 3d). While a diversity of timescales including long ones seem to be favorable for accurate modeling (Figure 3d and 3f), using constant memory timescales around 25 ms performs surprisingly well (matching the typical membrane timescales used in computational modeling Dayan and Abbott (2005), Figure 3e). Removing the hidden layer or decreasing the integration mechanism complexity deteriorates the performance significantly (Figure 3b). Allowing for more rapid changes of the memory through larger $\lambda$ is necessary (Figure 3c), possibly to match the fast internal dynamics of neurons around spike times.
@@
Within the ELM architecture, we allow for nonlinear interaction between any two synaptic inputs via the MLP. This flexibility might be necessary in cases where little is a priori known about the structure of the input. However, for matching the I/O of biological neurons, knowledge of neuronal morphology and biophysical assumptions about linear-nonlinear computations in the dendritic tree might be exploited to reduce the dimensionality of the input to the MLP, which so far is particularly costly due to the large number of synaptic inputs it receives ($d_{s}=1278$). In particular, many studies suggest individual dendritic branches compute a rectified sum of their synaptic inputs Poirazi et al. (2003); Stuart and Spruston (2015); Hawkins and Ahmad (2016); Poirazi and Papoutsi (2020), with additional nonlinearities located downstream. Consequently, we modify the ELM neuron to include virtual branches along which the synaptic input is first reduced by a simple summation before further processing (see Figure 4. For NeuronIO specifically, we assign the synaptic inputs to the branches in a moving window fashion (exploiting that in the dataset, neighboring inputs were also typically neighboring synaptic contacts on the same dendritic branch of the biophysical model). The window size is controlled by the branch size, and number of branches implicitly defines the stride size to ensure equally spaced sampling across the $d_{s}=1278$ inputs.
@@
Figure 4: Coarse-grained modeling of synaptic integration significantly improves model efficiency. a) The integration mechanism dynamics of the ELM now computes the activity of individual dendritic branches as a simple sum of their respective synaptic inputs first before passing them on to the $\text{MLP}_{{\color[rgb]{0.87109375,0.5625,0.01953125}\boldsymbol{w_{p}}}}$, where $d_{\mathrm{tree}}$ is the number of branches and $d_{\mathrm{brch}}$ the number of synapses per branch. b-c) Produced using joint ablation of $d_{m}$ with $d_{\mathrm{mlp}}=2d_{m}$. The new (branch) ELM neuron further improves on the vanilla ELM neuron by about 7$\times$ in terms of parameter efficacy. A (branch) ELM model with a humble 8104 trainable parameters now reaches 0.9924 AUC (3rd from left), using otherwise same hyper-parameters as the ELM (6th from right).
@@
Surprisingly, even with this strong simplification, the (branch) ELM neuron model is able to retain its predictive performance while requiring a modest 8K trainable parameters to substantially cross the performance threshold. This corresponds to roughly 7$\times$ reduction over the vanilla ELM neuron. We also find that a combination of $d_{\mathrm{tree}}=35$ and $d_{\mathrm{brch}}=65$ still achieved over $0.991$ AUC with only 5479 trainable parameters, corroborating the assumption of the near-linear computation within dendritic branches and inviting future investigation of minimal required synaptic nonlinearity. However, this simplification ideally utilizes the knowledge of the exact morphology for modeling the neuron, or the relationship structure in the input data in case of other tasks; thus most consecutive experimental reports are for the vanilla ELM neuron.
4.2 Evaluating temporal information integration capabilities on a bio-inspired task
@@
The Spiking Heidelberg Digits (SHD) dataset comprises spike-encoded spoken digits (0-9) in German and English Cramer et al. (2020b). The digits were encoded using 700 input channels in a biologically-inspired artificial cochlea. Each channel represents a narrow frequency band with the firing rate coding for the signal power in this band, resulting in an encoding that resembles the spectrogram of the spoken digit (see Figure 5a).
Motivated by recent findings that most neuromorphic benchmark datasets only require very limited temporal processing abilities Yang et al. (2021), we introduce the SHD-Adding dataset, created by concatenating two uniformly and independently sampled SHD digits and setting the target to their sum (regardless of language) (see Figure 5a). Solving this dataset necessitates identifying each digit on a shorter timescale and computing their sum by integrating this information over a longer timescale, which in turn requires retaining the first digit in memory. Whether single cortical neurons can solve this exact task is unclear; however, it has been shown that even single neurons possibly encode and perform basic arithmetics in the medial temporal lobe Cantlon and Brannon (2007); Kutter et al. (2018, 2022).
@@
Figure 5: The ELM neuron performs well on long and sparse data using longer timescales. a) Sample from the biologically motivated SHD-Adding dataset (based on Cramer et al. (2020b)), each dot is an input spike, and a vertical dashed line is a guide for the eye indicating the separation of the two digits (not communicated to the network). b-d) The ELM neuron consistently outperforms a classic LSTM, especially for smaller bin sizes (meaning longer training samples), and LSTM-performance cannot be fully recovered even for larger bin sizes. Our Leaky Integrate and Fire (LIF) Spiking Neural Network (SNN) does not manage to achieve good performance for any bin size, and training becomes unstable for long sequences. e) and f) Ablations using bin size of 2ms with test set performance reported. e) Solving SHD-Adding requires ELM neuron to have a higher complexity than required for NeuronIO. f) Longer memory timescales are crucial for extracting long-range dependencies in the data.
@@
The ELM neuron demonstrates its ability to solve the summing task across various temporal resolutions, as determined by the bin size. As we vary the bin size from 1ms (comprising 2000 bins in total, corresponding to the maximum temporal detail and longest required memory retention) to 100 (with 20 bins in total, corresponding to the minimum temporal detail and shortest memory retention), the ELM neuron’s performance remains robust, degrading only gracefully for greater bin sizes (see Figure 5b-d). Further, the performance is also maintained when testing on two held-out speakers, showing that the ELM neuron remains comparatively robust out-of-distribution. In contrast, the LSTM struggles with this task, especially when the bin size is below 50, due to memory issues. As the bin size increases, the LSTM’s performance does not improve because larger bin sizes leads to the loss of crucial temporal details. This outcome underlines the importance of a model’s ability to integrate complex synaptic information effectively (see Figure 5e). It also highlights the need for models to accommodate longer memory timescales, in light of the observation that limiting our model to memory timescales below 150 has a severe impact on performance (see Figure 5f).

### v3

@@ 2 The Expressive Leaky Memory Neuron
Spiking neurons emit their output spike at the axon hillock roughly when their membrane voltage crosses a threshold Kandel et al. (2000). The ELM neuron’s output is similarly based on its internal state ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{m_{t-1}}}$ (using a linear readout layer ${\color[rgb]{0.80078125,0.46875,0.73828125}\bm{w_{y}}}$), which rectified can be interpreted as the spike probability. For task compatibility, the output dimensionality is adjusted based on the respective dataset (not affecting neuron expressivity).
@@ 3 Related Work
Accurately replicating the full spike-level neuron input/output (I/O) relationship of detailed biophysical neuron models at millisecond resolution in a computationally efficient manner presents a formidable challenge. However, addressing this dynamics-learning task could yield valuable insights into neural mechanisms of expressivity, learning, and memory Durstewitz et al. (2023). The relative scarcity of prior work on this subject can be partially attributed to the computational complexity of cortical neurons only recently garnering increased attention Tzilivaki et al. (2019); Beniaguev et al. (2021); Larkum (2022); Poirazi & Papoutsi (2020). Additionally, traditional phenomenological neuron models have primarily aimed to replicate specific computational phenomena of neurons or networks Koch (1997); Izhikevich (2004); Herz et al. (2006), rather than the entire I/O relationship.
@@
Phenomenological neuron modeling research on temporally or spatially less detailed I/O relationship of biophysical neurons has been primarily centered around the use of multi-layered ANN structures in analogy to the neurons dendritic tree Poirazi et al. (2003); Tzilivaki et al. (2019); Ujfalussy et al. (2018). Similarly, we parametrize the synaptic integration with an MLP, while crucially extending this modeling perspective in several ways. Drawing upon the principles of classical phenomenological modeling via differential equations Izhikevich (2004); Dayan & Abbott (2005), our approach embraces the recurrent nature of neurons. We further consider the significance of hidden states ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{m}}$ beyond membrane voltage, as seen in prior works with predetermined variables Brette & Gerstner (2005); Gerstner et al. (2014). This addition enables us to flexibly investigate internal memory timescales ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\tau_{m}}}$.
@@
Deep learning architectures for long sequence modeling have seen a shift towards the explicit incorporation of timescales for improved temporal processing, as observed in recent advancements in RNNs, transformers, and state-space models Gu et al. (2021); Mahto et al. (2021); Smith et al. (2023); Ma et al. (2023). Such an explicit approach can be traced back to Leaky-RNNs Mozer (1991); Jaeger (2002); Kusupati et al. (2018); Tallec & Ollivier (2018), which use a convex combination of old memory and updates, as done in ELM using ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\kappa_{m}}}$. Whereas the classic time-varying memory decay mediated by implicit timescales Tallec & Ollivier (2018), is known from classic gated RNNs like LSTM Hochreiter & Schmidhuber (1997) and GRU Cho et al. (2014). In contrast to complex gating mechanisms, time-varying implicit timescales, or sophisticated large multi-staged architectures, the ELM features a much simpler recurrent cell architecture only using constant explicit (trainable) timescales ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\tau_{m}}}$ for gating, putting the major emphasis on the input integration dynamics using a single powerful MLP.
@@ 4 Experiments
In the experimental section of this work, we address three primary research questions. First, can the ELM neuron accurately fit a high-fidelity biophysical simulation with a small number of parameters? We detail this investigation in Section 4.1. Second, how can the ELM neuron effectively integrate non-trivial temporal information? We explore this issue in Section 4.2. Third, what are the computational limits of the ELM design? Discussed in Section 4.3. For training details, hyper-parameters, and tuning recommendations, please refer to the Appendix Section B and Table S1.
4.1 Fitting a complex biophysical cortical neuron model’s I/O relationship
@@
The NeuronIO dataset primarily consists of simulated input-output (I/O) data for a complex biophysical layer 5 cortical pyramidal neuron model Hay et al. (2011). Input data features biologically inspired spiking patterns (1278 pre-synaptic spike channels featuring -1,1 or 0 as input), while output data comprises the model’s somatic membrane voltage and output spikes (see Figure 2a and 2b). The dataset and related code are publicly available Beniaguev et al. (2021), and the models were trained using Binary Cross Entropy (BCE) for spike prediction and Mean Squared Error (MSE) for somatic voltage prediction, with equal weighting.
@@
Figure 2: The ELM neuron is a computationally efficient model of cortical neuron. a) detailed biophysical model of a layer 5 cortical pyramidal cell was used to generate the NeuronIO dataset consisting of input spikes and output spikes and voltage. b) and c) Voltage and spike prediction performance of the respective surrogate models, produced using joint ablation of $d_{m}$ with $d_{\mathrm{mlp}}=2d_{m}$ for ELM models. Previously around 10M parameters were required to make accurate spike predictions using a TCN Beniaguev et al. (2021), an LSTM baseline is able to do it with 266K, and our ELM and Branch-ELM neuron model require 53K and 8K respectively (3rd from left each), simultaneously achieving much better voltage prediction performance than the TCN. For comparison in terms of TP/FP Rate performance or FLOPS cost see Fig. 4c or S1 respectively. Additional comparisons to other phenomenological neuron models, such as LIF and ALIF, are provided in Table S4.
@@
Our ELM neuron achieves better prediction of voltage and spikes than previously used architectures for any given number of trainable parameters (and compute). In particular, it crosses the “sufficiently good” spike prediction performance threshold (0.991 AUC) as proposed in (Beniaguev et al., 2021) by using 50K trainable parameters, which is around 200$\times$ improvement compared to the previous attempt (TCN) that required around 10M trainable parameters, and 6$\times$ improvement over a LSTM baseline which requires around 266K parameters (see Figure 2c-d). Overall, this result indicates that recurrent computation is an appropriate inductive bias for modeling cortical neurons.
@@
We use the fitted model to investigate how many memory units and which timescales are needed to match the neuron closely. We find that around 20 memory units are required (Figure 3a) with timescales that are allowed to reach at least 25 ms (Figure 3d). While a diversity of timescales, including long ones, seems to be favorable for accurate modeling (Figure 3d and 3f), ELM with constant memory timescales around 25 ms performs sufficiently well (matching the typical membrane timescales in computational modeling Dayan & Abbott (2005), Figure 3e). Removing the hidden layer or decreasing the integration mechanism complexity significantly reduces performance (Figure 3b). Allowing for more rapid memory updates through larger $\lambda$ is crucial (Figure 3c), possibly to match the fast internal dynamics of neurons around spike times or to absorb information faster into memory (more details in Appendix A). When fitting the simple leaky-integrate-and-fire (LIF) or adaptive LIF, we reach a better prediction with only a few memory units (Figure S7).
Figure 3: The ELM neuron gives relevant neuroscientific insights. Ablations on NeuronIO of different hyperparameters of an ELM neuron with AUC $\approx 0.992$, and a Branch-ELM with the same default hyperparameters. The number of removed divergent runs marked with $1^{*}$. a) We find between 10 and 20 memory-like hidden states to be required for accurate predictions, much more than typical phenomenological models use Izhikevich (2004); Dayan & Abbott (2005). b) Highly nonlinear integration of synaptic input is required, in line with recent neuroscientific findings Stuart & Spruston (2015); Jones & Kording (2022); Larkum (2022). c) Allowing greater updates to the memory units is beneficial (see Appendix A). d-f) Ablations of memory timescale (initialization and bounding) range or (constant) value, with the default range being 1ms-150ms. Timescales around 25 ms seems to be the most useful (matching the typical membrane timescale in the cortex Dayan & Abbott (2005)); however, a lack can be partially compensated by longer timescales. g) and h) Ablating the number of branches $d_{\mathrm{tree}}$ and number of synapses per branch $d_{\mathrm{brch}}$ of the Branch-ELM neuron.
@@
Within the ELM architecture, we allow for nonlinear interaction between any two synaptic inputs via the MLP. This flexibility might be necessary in cases where little is a priori known about the input structure. However, for matching the I/O of cortical neurons, knowledge of neuronal morphology and biophysical assumptions about linear-nonlinear computations in the dendritic tree might be exploited to reduce the dimensionality of the input to the MLP (parameter-costly component with $d_{s}=1278$ inputs). Consequently, we modify the ELM neuron to include virtual branches along which the synaptic input is first reduced by a simple summation before further processing (see Figure 4). For NeuronIO specifically, we assign the synaptic inputs to the branches in a moving window fashion (exploiting that in the dataset, neighboring inputs were also typically neighboring synaptic contacts on the same dendritic branch of the biophysical model). The window size is controlled by the branch size $d_{\mathrm{brch}}$, and the stride size is derived from the number of branches $d_{\mathrm{tree}}$ to ensure equally spaced sampling across the $d_{s}=1278$ inputs.
@@
Figure 4: Coarse-grained modeling of synaptic integration significantly improves model efficiency. a) The integration mechanism dynamics of the ELM now computes the activity of individual dendritic branches as a simple sum of their respective synaptic inputs first before passing them on to the $\text{MLP}_{{\color[rgb]{0.87109375,0.5625,0.01953125}\bm{w_{p}}}}$, where $d_{\mathrm{tree}}$ is the number of branches and $d_{\mathrm{brch}}$ the number of synapses per branch. b) Accurate predictions using a Branch-ELM neuron with 8104 parameters (for zoomed-in version with model dynamics see Figure S9). c). The new Branch-ELM neuron improves on the ELM neuron by about 7$\times$ in terms of parameter efficacy (same ELM hyper-parameters). Differences in model quality are highlighted when examining a True-Positive rate at a low False-Positive rate.
@@
Surprisingly, even with this strong simplification, the Branch-ELM neuron model can retain its predictive performance while requiring 8K trainable parameters (roughly 7$\times$ reduction over the vanilla ELM) to cross the performance threshold substantially. We also find that a combination of $d_{\mathrm{tree}}=45$, $d_{\mathrm{brch}}=65$ and $d_{m}=15$ still achieved over $0.9915$ AUC with only 5329 trainable parameters, corroborating the assumption of the near-linear computation within dendritic branches and inviting future investigation of minimal required synaptic nonlinearity. However, this simplification utilizes the knowledge of morphology for modeling the neuron (in our case, exploiting the neighborhood in the dataset), violating it leads to deterioration of performance (Figure S6), therefore for most of the task we use the vanilla ELM neuron.
4.2 Evaluating temporal processing capabilities on a bio-inspired task
@@
The Spiking Heidelberg Digits (SHD) dataset comprises spike-encoded spoken digits (0-9) in German and English Cramer et al. (2020). The digits were encoded using 700 input channels in a biologically inspired artificial cochlea. Each channel represents a narrow frequency band with the firing rate coding for the signal power in this band, resulting in an encoding that resembles the spectrogram of the spoken digit (see Figure 5a).
@@
Figure 5: The ELM neuron performs well on long and sparse data using longer timescales. a) Sample from the biologically motivated SHD-Adding dataset (based on Cramer et al. (2020)), each dot is an input spike, and a vertical dashed line is a guide for the eye indicating the separation of the two digits (not communicated to the network). b-d) The ELM neuron (186K params.) consistently outperforms a classic LSTM (956K params.), especially for smaller bin sizes (meaning longer training samples), and LSTM-performance cannot be fully recovered even for larger bin sizes. The Branch-ELM (67K params.) can retain performance for fine-grained binning at a much reduced model size. Our LIF neuron based Spiking Neural Network (SNN) (51K params.) does not manage to achieve good performance for any bin size, and training becomes unstable for long sequences. e) and f) Ablations using a bin size of 2ms with test set performance reported. e) Solving SHD-Adding requires ELM neuron to have a higher complexity than required for NeuronIO, and much larger models become unstable. Potentially a network of smaller ELM neuron might be preferable. f) Longer ${\color[rgb]{0.00390625,0.44921875,0.69921875}\tau_{m}}$ are crucial for extracting long-range dependencies. Possibly shorter ones might suffice in a ELM network, as longer timescales can emerge through dynamics Khajehabdollahi et al. (2023).
Motivated by recent findings that most neuromorphic benchmark datasets only require minimal temporal processing abilities Yang et al. (2021), we introduce the SHD-Adding dataset by concatenating two uniformly and independently sampled SHD digits and setting the target to their sum (regardless of language) (see Figure 5a). Solving this dataset necessitates identifying each digit on a shorter timescale and computing their sum by integrating this information over a longer timescale, which in turn requires retaining the first digit in memory. Whether single cortical neurons can solve this exact task is unclear; however, it has been shown that even single neurons possibly encode and perform basic arithmetics in the medial temporal lobe Cantlon & Brannon (2007); Kutter et al. (2018; 2022).
@@
The ELM neuron solves the summing task across various temporal resolutions (determined by the bin size). As we vary the bin size from 1ms (2000 bins in total, the maximal temporal detail and longest required memory retention) to 100 (20 bins in total, the minimal temporal detail and shortest memory retention), the ELM neuron’s performance remains robust, degrading proportionally to the bin sizes (see Figure 5b-d); this drop in performance is not a shortcoming of the model itself, but a consequence of loss of temporal information through binning. Further, the performance is also maintained when testing on two held-out speakers, showing that the ELM neuron remains comparatively robust out-of-distribution. Due to vanishing gradients, the LSTM performs worse on this task, especially when the bin size is below 50. As the bin size increases, the LSTM’s performance improves but does not surpass ELM because larger bin sizes likewise lead to the loss of crucial temporal details. This outcome underlines the importance of a model’s ability to integrate complex synaptic information effectively (see Figure 5e) and the utility of longer neuron-internal timescales for learning long-range dependencies, potentially necessary for cortical neuron’s operation (see Figure 5f).